Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image


Mission Operations Days: 2001/279 to 2001/285
October 06, 2001 00:00:00 GMT to October 12, 2001 00:00:00 GMT

NOTE: A new Table #5: Ideal Blackbody Output, was loaded into the instrument on Friday, October 12, 2001 (2001/285) at approximately 13:45:00 GMT. The following changes were made:

Lowered values for Bands 5 to 7 by 200 (decimal) counts.
Increased values for Bands 20, 22 to 25, 27 and 28 by 100 counts

If the new value would have been larger than 4095, then the value was set to 4095. This happened a couple of times for higher temperature values of B24, B25, B27, and B28.

Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On; #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 3 set of patches
  • FR Error Counting and Statistics (2 patches)
  • FR Band 5 Gain Change (20 commands)
  • FR BB Ideal Output Values (1 patch)
Inhibit Ids SetNoneNominal
TMON 66 (SRCA A)DisabledNominal
TMON 67 (SRCA B)DisabledNominal
TMON 68 (FIFO)EnabledNominal
TMON 69 (EPS Current)DisabledNominal

This Week's Completed MODIS Activities:

Saturday, October 6, 2001


Sunday, October 7, 2001

Lunar Roll Maneuver (-11.3408 degree roll)
2001/280 05:28:52 Real-time FOT sets IRU to high rate (lower spacecraft attitude precision)
2001/280 06:36:57 ATC Moved Formatter to Day Rate ahead by 03:23 to ensure 50/50 day/night data collection
2001/280 07:10:46 ATC Set formatter to Day Rate
2001/280 07:10:48 ATC Set SCIABNORM flag to ABNORM
2001/280 07:10:50 ATC Set PV DCR computation to OFF
2001/280 07:10:52 ATC Set PC DCR computation to OFF
2001/280 07:10:54 ATC Set Sector Rotation to –3072
2001/280 07:17:24 ATC Set Sector Rotation to zero
2001/280 07:17:26 ATC Set PC DCR computation to ON
2001/280 07:17:28 ATC Set PV DCR computation to ON
2001/280 07:17:30 ATC Set SCIABNORM flag to NORM
2001/280 07:17:32 ATC Set formatter to Night Rate
2001/280 07:33:09 ATC Moved Formatter to Night Rate later by 03:23 to ensure 50/50 day/night data collection
2001/280 07:39:44 Real-time FOT sets IRU to low rate (greater spacecraft attitude precision)

Monday, October 8, 2001


Tuesday, October 9, 2001


Wednesday, October 10, 2001

2001/283 18:25:03 – 18:30:30 Real-time Memory Dump of first half of table 5
2001/283 19:18:12 – 19:24:00 Real-time Memory Dump of second half of table 5
2001/283 21:01:28 ATC OA-15: SD/SDSM Open
2001/283 21:52:52 – 22:26:25 ATC OA-19: SRCA Full Radiometric
2001/283 22:40:21 ATC OA-16: SD/SDSM Screened

Thursday, October 11, 2001

2001/284 06:05:30 – 06:57:09 ATC OA-22: SRCA Full Spectral Part 1 of 4
2001/284 08:00:56 – 08:35:41 ATC OA-22: SRCA Full Spectral Part 2 of 4
2001/284 09:22:58 – 10:14:14 ATC OA-22: SRCA Full Spectral Part 3 of 4
2001/284 11:17:00 – 11:54:58 ATC OA-22: SRCA Full Spectral Part 4 of 4
2001/284 15:15:35 – 15:18:06 ATC OA-27: PV Ecal
2001/284 15:18:35 – 15:23:06 ATC OA-28: PC Ecal

Friday, October 12, 2001

2001/285 13:40:00 Real-time Set SCIABNORM flag to ABNORM
2001/285 13:41:33 – 13:51:15 Real-time Memory Load Table 5 New Ideal Blackbody output table, see changes at top of report
2001/285 13:53:23 Real-time Set PV and PC DC Restore to ON They were commanded to OFF during the table load
2001/285 13:54:24 Real-time Set SCIABNORM flag to ABNORM
2001/285 13:55:30 – 14:06:05 Real-time Memory Dump of Table 5

This Week's Scheduled MODIS Activities Not Completed:


Upcoming MODIS Events:

A blackbody calibration and two SD/SDSM calibrations will be scheduled during the upcoming week.


The next drag make-up maneuver is scheduled to occur on October 23, 2001 (2001/296).

The next Lunar Calibration roll maneuver is scheduled for day 2001/309 (November 5, 2001) with data collection from 18:15:11 to 18:21:41. The roll angle will be -11.8237 degrees.

MODIS Anomalies:

Formatter Error Events:

The FR error counter (MOD_SS_CP_STATUS_08) has “rolled over” 18 times since the A-Side restart following the PS-2 Anomaly, and 16 times since the latest string of errors started on August 15. This is a 16-bit signed counter, therefore it “rolls over” every 32,768 counts.

The error rate had been increasing (up to ~36,000 per day) until a series of instrument calibrations on Wednesday and Thursday of this week. Two SD calibrations and two SRCA calibrations heated up the instrument (MOD_TP_FR_A_ENGINE >2C) such that the formatter errors ceased for about seven hours. As the instrument cooled, errors resumed but at a severly diminished rate.

Two error tracking spreadsheets have been updated and can be retrieved from:

An MS Excel spreadsheet called, AM1_FR_error_rollovers_101201.xls updates the version from October 9th with the latest data available. A second spreadsheet, MOD_ME_Orbit_temps_2001_195-285.xls contains 90 days of orbital average temperatures for a wide variety of main electronics module mnemonics. Only one plot (of the FR A/B temperatures) has been included, but the data are available for other such plots.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

Note: SD/SDSM calibrations were planned to be reduced to once per two weeks. This plan has been revised; SD/SDSM calibrations will continue to be scheduled once each week until the end of November.

MODIS Telemetry Trends:

Instrument temperatures across the board are risisng slightly as of late, although the orbit beta angle has not yet reached its minimum.

Non-MODIS Significant Events:

The Network Communications Center (NCC) has been rejecting and/or cancelling Terra TDRSS contact times (likely to support DoD satellites). This has resulted in at least one incidence of data loss.

Limited Life Item Status:

SRCA 10W Lamp #1: 192.2 of 500 hours
SRCA 10W Lamp #2: 146.6 of 500 hours
SRCA 10W Lamp #3: 155.8 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 558.3 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1727 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements